Search This Blog


Saturday / Sunday 6 & 7 November 2010 ~ Out of the Picture

SATURDAY. I was absent from the area on Saturday until about 17.00. Fortunately, I had time to pop over to King's Lynn during the evening to catch up with events.
Assuming nothing had changed, it looks very much like today was a completely low floor service and we haven't had that for a long while. K18 had 37574 taken off this afternoon, but 37160 looks like it was the replacement.L12 had a choice from 37567, 37157 or 'The Beast' so we should find out eventually what was involved. Tonight 37575 on K15 was replaced for the last trip by 37563 which had done K14.
37569 was fitted with a new wheel hub trip meter today, the first B9 to receive a third such device.

SUNDAY. Not a lot to report today. 37564 is the latest B9 to suffer from the noisy fan fault. Recent examples have all been B7s and 564 is the first B9 for some time. Today it was replaced in service by 37568.

No comments: